OpenStudy (anonymous):

Write expression in radical form!? 1.(3n)^3/4 2.(5x)^7/4 3.(4b)2/3 4.x^7/4

jimthompson5910 (jim_thompson5910):

I'll do the first one to get you started

jimthompson5910 (jim_thompson5910):

Use the idea that \[\large x^{m/n} = \sqrt[n]{x^m}\] to get \[\large (3n)^{3/4} = \sqrt[4]{(3n)^3}\] \[\large (3n)^{3/4} = \sqrt[4]{3^3*n^3}\] \[\large (3n)^{3/4} = \sqrt[4]{27n^3}\]

OpenStudy (anonymous):

Alright, for the next one, (5x)^7/4, i got \[\sqrt[4]{78125x^7}\]

OpenStudy (anonymous):

That doesnt seem right..... @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's correct

OpenStudy (anonymous):

Serously? Sweet! Thanks for teaching me instead of just giving the answer! :)

jimthompson5910 (jim_thompson5910):

you're welcome